what is the solution of this problem 7-(4z-9)=9-5z please reply



























Answer by sofiyac(983)

7-(4z-9)=9-5z
7-4z+9=9-5z you can cancel the 9's
7-4z=-5z add 4z to each side
7=-z divide each side by -1
-7=z